# FuncGenFoil: Airfoil Generation and Editing Model in Function Space

### v1

@@ 4.3 Ablation Study
We first ablate the use of the Matérn kernel $K$ with different parameters as the kernel function for the sampled Gaussian process. As shown in Table 4, we compare the cases of Matérn kernel $\nu=1.5$, $l=0.01$, and Matérn kernel $\nu=2.5$, $l=0.03$, where the latter is smoother due to the combined effect of both parameters. Notably, the limits of $l=0$ and $\nu\to\infty$ correspond to pure white noise and the RBF kernel, respectively. To further highlight the importance of kernel function selection, we also test two models of different sizes (by setting different FNO modes). From Table 4, we observe that a smoother kernel improves both the smoothness and accuracy of the generated airfoils. Additionally, the comparison highlights that, under our setup, selecting a smoother kernel leads to a more substantial performance improvement than increasing the model size. This underscores the benefit of modeling the function object in the airfoil generation task rather than modeling discrete points.
In Table 5, we compare the MSE of constraint condition calculations in the airfoil editing task using two initialization schemes: (1) using the latent variables obtained through ODE inversion of the original airfoil as the prior, and (2) using a zero prior $u_{0}=0$ without ODE inversion. Both schemes are fine-tuned with the same number of iterations. As shown, the ODE inversion-based prior significantly reduces the MSE compared to the zero prior, with a maximum reduction of 84.85. This underscores the importance of incorporating original airfoil information during the initialization process.
Table 4: Ablation study on the impact of different kernels and model sizes for the conditional airfoil generation task.
Matérn $\nu$ Matérn $l$ FNO modes $\bar{\sigma}\downarrow$ $\mathcal{M}\downarrow$
1.5 0.01 32 0.108 0.0119
2.5 0.03 32 0.013 0.0098
1.5 0.01 48 0.039 0.0101
2.5 0.03 48 0.013 0.0098
@@ 5 Conclusion
Appendix A Airfoil Geometry Parameters
Table 6: Geometric parameters of the airfoils used in conditional generation tasks
Index Symbol Meaning
1 $R_{\mathrm{le}}$ the leading edge radius
2 $X_{\mathrm{up}}$ upper crest position x
3 $Y_{\mathrm{up}}$ upper crest position y
4 $Z_{\mathrm{xxup}}$ upper crest curvature
5 $X_{\mathrm{lo}}$ lower crest position x
6 $Y_{\mathrm{lo}}$ lower crest position y
7 $Z_{\mathrm{xxlo}}$ lower crest curvature
8 $Y_{\mathrm{te}}$ trailing edge position
9 $\Delta Y_{\mathrm{te}}$ trailing thickness
@@ 11 $\beta_{\mathrm{te}}$ trailing edge angle down
Algorithm 2 Generative Model Training.
@@
Algorithm 3 Inference Process (Airfoil Generation)
@@
Table 7: Training hyperparameters and model parameters in conditional airfoils generation tasks
@@
max learning rate 5e-6
batch size 512
optimizer used in pre-training and finetuning Adam
train max iterations 1,000,000
ODE Solver time steps 100
@@
Fourier neural operator layers 8
Fourier neural operator modes 48
@@
Matérn kernel $\nu$ 2.5
Matérn kernel $l$ 0.03
We adopted the same conditions and symbols as those in AFBench. The parameters below, including the leading-edge radius, the curvature at the maximum thickness of the upper and lower surfaces, and the trailing-edge angle of the airfoil, were obtained through fitting. Note that our implementation of the angle calculation differs slightly from that of AFBench, as we found that the original fitting method was not invariant when the sampling precision increased.
Appendix B Algorithm Details
The algorithm for training airfoil generative models is shown in Algorithm 2, and inference using a trained model is shown in Algorithm 3. In the actual implementation, we utilized Optimal Transport techniques and employed the Matérn kernel as the kernel function for the Gaussian process in the latent space of the generative model.
@@ Appendix C Training Details
In most of the experiments, we followed the hyperparameters listed in the table below to train the models, including both the pre-training and fine-tuning stages. On the AF-200K dataset, we increased the maximum number of iterations to 2,000,000 and the batch size to 2048. For the number of time steps used to solve the ODE, we tested reconstruction metrics on the super dataset with 10, 100, and 1000 steps, and ultimately chose 100 steps as the default setting for all other experiments, except for the fine-tuning stage, where we used 10 steps for training efficiency.

### v3

@@ 4.3 Ablation Study
Figure 6: (A): Comparison of the geometric mean error $\bar{\sigma}_{g}$ using kernels of different smoothness but same characteristic length $l=0.03$. (B): Comparison of the geometric mean error $\bar{\sigma}_{g}$ using Matérn kernel $\nu=3.5$ with different characteristic length $l$. (C): Comparison of the geometric mean error $\bar{\sigma}_{g}$ using different Fourier mode in FNO using Matérn kernel of same $l=0.03$.
Kernels of Gaussian Process. The type of Gaussian process kernel used as a prior in the model influences the function space into which the data is diffused. We perform ablation experiments across several kernel types, including the white noise kernel, the Matérn kernel with smoothness parameter $\nu\in[1.5,2.5,3.5]$, and the radial basis function (RBF) kernel. We measure the geometric mean $\bar{\sigma}_{g}$ under different smoothness, characteristic length, and Fourier mode settings, as shown in Figure 6. We observe that FunGenfoil benefits from employing smoother kernels, which enhance both the smoothness and accuracy of the generated airfoils. This effect arises because the parameter $\nu$ primarily controls the smoothness of the function space, and the RBF kernel corresponds to the case $\nu\rightarrow\infty$. Additionally, we find that there exists an optimal characteristic length at $l=0.03$ for the kernel. In our experiments, FunGenfoil does not exhibit improved performance with an increased number of Fourier modes. This phenomenon may be due to the predominantly low-frequency characteristics of supercritical airfoils, such that adding higher modes in the FNO layer introduces unnecessary model complexity. More detailed experimental results are provided in Table 8 in the Appendix.
@@
Editing with Prior via ODE Inversion. In Table 4, we compare MSE of constraint condition calculations in the airfoil editing task using two initialization schemes: (1) using the latent variables obtained through ODE inversion of the original airfoil as the prior, and (2) using a zero prior $u_{0}=0$ without ODE inversion. Both schemes are fine-tuned with the same number of iterations. The ODE inversion-based prior significantly reduces the MSE compared to the zero prior, with a maximum reduction of 84.85. This underscores the importance of incorporating original airfoil information during the initialization process.
ODE Numerical Method. Model inference involves solving an ordinary differential equation (ODE), and the choice of numerical integration scheme affects its performance. Using the same velocity operator $v_{\theta}$ trained on the supercritical airfoil dataset and sampling from a Gaussian prior with a Matérn kernel ($\nu=2.5$, $l=0.03$), we evaluate FuncGenFoil across various time step sizes and ODE solvers (Euler, midpoint, and fourth-order Runge-Kutta (RK4) methods). Results in Table 5 show that FuncGenFoil benefits from larger time steps using Euler method. However, RK4 may degrade performance at excessively small step sizes, potentially because the model enters regions where the velocity operator is inadequately trained.
Table 5: Quantitative experiments on the impact of ODE solver and time step on sampling quality.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
Euler $10$ 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
$50$ 0.62 3.61 0.07 92.5 2.08 0.06 96.8 0.06 0.07 14.5 15.7 20.6 1.60 -107.6 0.99
$100$ 0.67 3.13 0.09 83.9 1.84 0.08 90.6 0.07 0.06 12.3 13.1 18.7 1.55 -106.4 0.99
Midpoint $10$ 0.68 5.29 0.08 161.3 3.12 0.08 178.9 0.06 0.09 21.4 22.4 35.8 2.17 -105.1 1.00
$50$ 0.75 3.82 0.12 137.9 2.34 0.10 153.4 0.08 0.06 14.1 14.9 29.8 1.97 -103.3 0.99
$100$ 0.82 3.94 0.14 125.7 2.22 0.13 137.4 0.08 0.07 14.4 14.2 27.2 2.06 -101.6 0.99
RK4 $10$ 15.9 46.3 1.21 4406 25.7 1.24 4795 1.17 1.68 333 396 911 36.6 -67.5 3.05
$50$ 0.92 7.63 0.24 482.7 5.55 0.22 511.8 0.16 0.19 39.3 41.3 99.1 4.68 -92.9 1.07
$100$ 0.82 5.65 0.18 160.6 3.21 0.17 170.8 0.10 0.12 22.0 22.8 35.1 2.78 -96.9 1.00
@@ 5 Conclusion
Appendix A Problem Setting for Airfoil Generation
A.1 Airfoil Parametrization
In this paper, we perform a circular parameterization of the airfoil curves, as illustrated in Figure 7. Specifically, we project a unit circle onto the $x$-axis and adopt the angle $\alpha$ as the function domain. The parameterized airfoil function begins at the trailing edge on the suction side, where $\alpha=0$, and ends at the trailing edge on the pressure side, where $\alpha=1$. In this manner, we transform the airfoil data from a circular topology into a well-defined functional form suitable for reconstruction by generative models.
Figure 7: Illustration of the airfoil curve parameterization employed in FuncGenFoil.
A.2 Design Parameters of Airfoil Geometry
Index Symbol Geometric Meaning
1 $\mathrm{R_{le}}$ leading edge radius
2 $\mathrm{X_{up}}$ upper crest position x
3 $\mathrm{Y_{up}}$ upper crest position y
4 $\mathrm{Z_{xxup}}$ upper crest curvature
5 $\mathrm{X_{lo}}$ lower crest position x
6 $\mathrm{Y_{lo}}$ lower crest position y
7 $\mathrm{Z_{xxlo}}$ lower crest curvature
8 $\mathrm{Y_{te}}$ trailing edge position
9 $\Delta\mathrm{Y_{te}}$ trailing thickness
@@ 11 $\beta_{\mathrm{te}}$ trailing edge angle down
(a) PARSEC parameters and its geometric meaning.
(b) Demonstration of PARSEC parameters.
Table 6: Eleven PARSEC parameterization in airfoil design.
Algorithm 1 Model Training.
@@
Algorithm 2 Model Inference.
@@
Algorithm 3 Model Finetuning (Airfoil Editing).
Output: new airfoil $\{y_{i}\}$ at resolution $d$.
2:  Set $\{a_{\theta,i}\}=\{u_{0,i}\}$.
3:  while finetuning… do
4:     Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
[…]
6:     Compute gradient and update $\theta$.
7:  end while
8:  Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
9:  return $\{y_{i}\}=\{u_{1,i}^{{}^{\prime}}\}$
Table 7: Training hyperparameters and model parameters in conditional airfoil generation tasks
@@
Max learning rate $5\times 10^{-6}$
Batch size $1024$
Batch size (AF-200K) $2048$
Optimizer (pre-training and fine-tuning) Adam
Optimizer scheduler Cosine Annealing
Warmup iterations $2000$
Max training iterations $1000000$
Max training iterations (AF-200K) $2000000$
ODE solver time steps $10$
ODE solver type Euler
@@
Fourier neural operator layers $6$
Fourier neural operator modes $64$
@@
Gaussian process kernel Matérn
Kernel noise scale $n$ $1$
Kernel order $\nu$ $2.5$
Kernel characteristic length $l$ $0.03$
[…]
Appendix B Algorithm Implementation Details
The algorithm for training the airfoil generative model is presented in Algorithm 1, and the inference procedure using a trained model is described in Algorithm 2. In our implementation, we employed Optimal Flow Matching and utilized a Gaussian process with specific kernel functions, such as the RBF kernel or the Matérn kernel, in the latent space of the generative model.
@@ Appendix C Training Details
In most of the experiments, we followed the hyperparameters listed in the Table 7 below to train the models, including both the pre-training and fine-tuning stages. On the AF-200K dataset, we increased the maximum number of iterations to $2000000$ and the batch size to $2048$. For the number of time steps used to solve the ODE, we use $10$ steps for relative balance between fast sampling speed and good performance, except for the fine-tuning stage, where we used 10 steps for training efficiency.
Appendix D More Experiments
We conduct a comprehensive ablation study on the choice of Gaussian process kernel types, model modes of Fourier Neural Operators, and characteristic lengths, as shown in Table 8 for the Matérn kernel, Table 9 for the radial basis function (RBF) kernel, and Table 10 for the white noise kernel, which is commonly used in diffusion or flow models in finite-dimensional spaces. In general, appropriate selections of these settings can significantly enhance model performance. Specifically, the Gaussian prior should exclude function spaces where the data is unlikely to reside, making the prior space as restricted as possible to facilitate effective model training. For example, a dataset with high smoothness is unlikely to belong to the function space generated by a white noise prior but is more plausibly sampled from a Gaussian prior with an RBF kernel.
In Table 9, we observe that the RBF kernel exhibits a similar tendency to the Matérn kernel, with both achieving optimal performance at the characteristic length $0.03$. As for the white noise kernel in Table 10, since its characteristic length is effectively $0$, we studied the influence of the noise scale $n$ over the range $[1.0,2.0,4.0]$. We found that excessively large diffusion white noise negatively impacts model performance, likely because such noise scales project the function into a larger white noise space, making it more difficult for the neural operator to learn the intrinsic dynamics.
Table 8: FuncGenFoil model performance with different FNO mode and Matérn kernel for conditional airfoil generation.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
$32$ $2.5$ $0.01$ 0.62 6.06 $\mathbf{0.09}$ 200.4 3.58 0.09 185.3 $\mathbf{0.07}$ 0.12 26.7 29.1 41.1 2.51 -105.6 1.02
$0.02$ 0.52 5.10 0.10 97.4 3.26 0.11 107.0 0.08 0.12 26.2 $\mathbf{27.6}$ 24.3 2.22 -105.7 $\mathbf{0.99}$
$0.03$ 0.54 $\mathbf{4.78}$ 0.11 $\mathbf{82.3}$ 3.00 0.11 93.2 0.09 0.12 $\mathbf{26.0}$ 27.7 $\mathbf{21.6}$ $\mathbf{2.18}$105.6 $\mathbf{0.99}$
$3.5$ $0.01$ 0.63 5.64 $\mathbf{0.09}$ 158.3 3.46 $\mathbf{0.08}$ 156.9 $\mathbf{0.07}$ 0.12 27.6 29.2 34.7 2.41 -105.8 1.01
$0.02$ $\mathbf{0.48}$ 5.08 0.10 89.4 3.12 0.10 97.5 0.08 0.12 26.8 28.2 22.8 $\mathbf{2.18}$105.9 $\mathbf{0.99}$
$0.03$ 0.53 5.17 0.11 83.5 $\mathbf{2.90}$ 0.11 $\mathbf{89.5}$ 0.10 0.14 29.4 30.7 22.0 2.28 -105.2 $\mathbf{0.99}$
$64$ $1.5$ $0.01$ 0.82 8.57 0.10 370.2 5.21 0.09 342.5 0.08 0.12 36.3 37.5 72.9 3.30 -104.0 1.09
$0.02$ 0.65 8.26 0.13 222.9 4.97 0.11 219.6 0.08 0.13 37.4 35.9 48.2 3.12 -103.6 1.04
$0.03$ 0.71 8.18 0.14 183.7 4.82 0.12 207.2 0.09 0.14 34.6 36.6 43.3 3.13 -104.0 1.02
$2.5$ $0.01$ 0.80 8.77 0.10 291.2 5.23 0.10 277.6 0.08 0.12 35.8 37.7 59.8 3.20 -103.7 1.06
$0.02$ 0.68 8.66 0.13 192.4 4.78 0.11 182.8 0.09 0.12 36.2 37.8 42.2 3.06 -103.5 1.02
$0.03$ 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
$3.5$ $0.01$ 0.88 8.80 0.11 269.3 5.26 0.09 248.9 0.08 0.12 36.5 38.0 55.3 3.20 -103.4 1.05
$0.02$ 0.72 8.52 0.11 177.1 4.84 0.11 174.6 0.08 0.13 33.2 37.4 39.7 2.98 -103.6 1.02
$0.03$ 0.66 7.80 0.12 161.1 4.54 0.11 164.2 0.09 0.15 35.1 38.3 37.5 2.96 -103.4 1.01
$0.04$ 0.82 8.19 0.13 170.0 4.45 0.13 159.7 0.09 0.16 36.6 34.8 37.7 3.11 -103.6 1.01
$0.06$ 0.87 8.82 0.16 183.0 5.01 0.13 183.9 0.10 0.17 41.8 41.2 42.3 3.44 -102.8 1.02
$0.12$ 1.39 10.7 0.20 261.0 6.23 0.15 230.3 0.12 0.19 53.4 44.9 55.3 4.27 $\mathbf{-99.9}$ 1.05
$128$ $2.5$ $0.01$ 1.09 10.4 0.11 708.5 7.13 0.10 613.5 0.09 $\mathbf{0.11}$ 40.7 41.1 129.3 4.16 -102.8 1.27
$0.02$ 1.15 10.0 0.11 577.5 6.83 0.11 570.2 0.09 0.14 37.8 42.3 113.3 4.14 -103.4 1.20
$0.03$ 1.05 10.1 0.12 598.0 6.48 0.14 487.7 0.09 0.15 36.9 40.1 107.3 4.22 -103.0 1.23
$3.5$ $0.01$ 1.04 10.3 0.10 628.9 7.05 0.09 576.8 0.08 0.12 39.3 42.0 118.7 4.05 -102.9 1.23
$0.02$ 0.93 9.94 0.12 560.6 6.81 0.10 484.6 0.09 0.15 41.5 40.8 104.1 4.05 -103.2 1.19
$0.03$ 1.11 9.93 0.11 548.2 6.49 0.12 520.3 0.09 0.16 38.6 39.0 105.8 4.16 -103.1 1.23
Table 9: FuncGenFoil model performance with different RBF kernel for conditional airfoil generation.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
$64$ $0.001$ 1.26 $\mathbf{7.60}$ $\mathbf{0.09}$ 1224 5.34 $\mathbf{0.07}$ 1046 $\mathbf{0.06}$ $\mathbf{0.08}$ 24.9 $\mathbf{28.6}$ 213 3.68 -103.7 1.42
$0.01$ 0.77 8.67 0.12 223.6 5.05 0.09 174.7 0.08 0.12 35.2 37.1 47.0 3.06 -103.6 1.03
$0.02$ $\mathbf{0.62}$ 8.33 0.12 162.2 4.59 0.11 165.9 0.09 0.17 38.0 42.5 38.4 3.04 -103.2 $\mathbf{1.01}$
$0.03$ 0.66 7.80 0.12 $\mathbf{147.9}$ $\mathbf{4.43}$ 0.12 $\mathbf{154.2}$ 0.09 0.15 $\mathbf{33.2}$ 37.9 $\mathbf{35.1}$ $\mathbf{2.93}$103.6 1.00
$0.04$ 0.84 8.79 0.13 197.1 4.59 0.12 168.4 0.09 0.17 35.8 37.9 41.3 3.20 -103.2 1.02
$0.06$ 1.15 9.97 0.15 220.3 6.21 0.14 215.5 0.10 0.16 38.3 44.2 48.7 3.73 -99.9 1.05
$0.12$ 1.12 10.9 0.17 230.5 7.71 0.17 265.6 0.11 0.18 40.0 42.3 54.4 4.14 $\mathbf{-95.1}$ 1.08
Table 10: FuncGenFoil model performance with white noise kernel of different noise scale $n$ for conditional airfoil generation.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
$64$ 1 1.15 7.18 0.09 913.8 5.16 0.07 964.5 0.06 0.08 26.2 26.1 176.8 3.44 -104.7 1.39
2 1.46 7.98 0.10 1274 5.94 0.08 1268 0.08 0.10 30.6 32.4 238.3 4.17 -102.7 1.64
4 2.27 9.14 0.12 2015 7.24 0.11 2059 0.09 0.14 41.5 41.4 379.7 5.61 -99.9 2.09
Appendix E Limitations
Although we have taken a step toward general object shape modeling in function space, the focus of this paper is primarily on airfoils or aircraft wings, which have relatively simple geometries and smooth surfaces yet are highly significant for aerodynamic performance. Thus, our current approach has limitations regarding the scope of geometry. In future studies, if we aim to extend our method to modeling entire aircraft or dealing with objects of general shapes, substantial theoretical work and experimental analysis will be required, especially when a suitable coordinate system is unavailable for describing complex, irregular, or non-smooth geometries. This direction is part of our ongoing efforts.
Appendix F Societal Impacts
This work will accelerate research in AI for Science and Engineering and generally has positive social impacts. On the other hand, AI models for engineering design and optimization could accelerate technological diffusion within society, potentially raising issues related to intellectual property or unauthorized technology transfer to entities or individuals intending to design specific engineering products or construct items harmful to society.